For the numbers below write their prime factorization and compute their gcd and lcm:
(i) 831600 (ii) 70560

Answer by tommyt3rd(5050)

831600 =2^4*3^3*5^2*7*11
70560 = 2^5*3^2*5*7^2
gcf(831600,70560)=2^4*3^2*5*7 = 5040
lcm(831600,70560)=2^5*3^3*5^2*7^2*11 = 11642400

:)